# PAOLI: Pose-free Articulated Object Learning from Sparse-view Images

### v1

@@ 3.4 Extension to Multiple Moving Parts
For simplicity, we have described our method for two parts, but it can be easily extended to objects with multiple parts as shown in Fig. 3. This extension requires additional input images for the new moving part which are used to initialize the 3D Gaussian splats and camera parameters for each set of images in §3.1.Here we use the Gaussian splats from the second row of input images as source Gaussian and the first row as target Gaussian shown in Fig. 3. We first go through the pipeline with Gaussian deformation in §3.2 and ¢3.3 to model the motion of the first part, which is the light blue arm shown in Fig. 3. After the optimization, we obtain the articulation parameters and segmentation of the first arm. Based on this, we use the same source Gaussian together with the Gaussian in the bottom row as target Gaussian to model the second part by repeating deformation §3.2 and joint optimization §3.3. Note that there is always only 1 articulated part between source and target Gaussian.
@@ 4 Experiments
[…]
Motion Ang Err $\downarrow$ AGS[10] GT 62.083 36.568 86.058 78.129 62.083 54.613 88.677 87.454 59.562 75.744
AYN[7] GT 52.986 60.834 2.105 42.101 81.157 88.316 74.938 24.850 failed 66.718
AGS[10] VGGT 49.001 66.191 82.359 51.069 21.378 40.371 288.473 83.680 67.020 71.801
Ours 0.095 0.914 1.020 2.365 0.839 1.337 2.905 1.737 0.878 0.525
Pos Err $\downarrow$ AGS[10] GT 0.032 0.124 1.805 0.252 0.032 0.149 1.209 0.231 - -
AYN[7] GT 0.021 0.073 0.033 2.116 0.221 0.231 0.916 0.334 - -
AGS[10] VGGT 0.455 0.254 0.033 0.775 0.741 0.090 0.412 0.513 - -
Ours 0.180 0.007 0.106 0.042 0.058 0.196 1.042 0.157 - -
Motion Dist $\downarrow$ AGS[10] GT 70.900 75.548 46.925 29.381 70.900 64.865 61.955 33.345 1.275 0.213
AYN[7] GT 122.730 562.642 10.115 20.583 111.252 118.564 50.044 81.205 failed 0.373
AGS[10] VGGT 104.417 54.799 43.799 45.248 27.312 57.388 53.142 34.177 0.202 0.297
Ours 0.861 2.279 0.788 2.654 1.516 0.930 2.227 2.926 0.017 0.118
Rendering PSNR $\uparrow$ AGS[10] GT 9.349 11.294 13.787 13.871 13.481 12.898 15.856 17.644 24.543 14.333
AYN[7] GT 10.827 14.631 13.411 15.540 14.675 15.826 16.801 15.784 failed 14.018
AGS[10] VGGT 10.740 16.403 17.737 15.123 18.755 14.866 12.990 16.082 18.268 18.483
Ours 24.659 26.920 26.491 23.434 28.823 27.900 27.016 25.408 29.041 25.493
SSIM $\uparrow$ AGS[10] GT 0.652 0.722 0.809 0.723 0.846 0.831 0.842 0.877 0.962 0.765
AYN[7] GT 0.718 0.893 0.852 0.816 0.904 0.881 0.855 0.889 failed 0.814
AGS[10] VGGT 0.693 0.861 0.883 0.801 0.910 0.890 0.821 0.878 0.905 0.874
Ours 0.940 0.963 0.955 0.931 0.967 0.953 0.956 0.966 0.974 0.927
Geometry CD $\downarrow$ AGS[10] GT 1188.824 717.780 893.149 244.862 1188.824 1599.261 467.633 268.377 90.288 231.355
AYN[7] GT 677.761 164.321 124.728 335.108 258.443 125.208 740.898 1575.325 failed 62.890
AGS[10] VGGT 476.935 48.592 77.646 26.251 21.667 112.274 632.828 24.736 26.769 15.599
Ours 0.856 2.051 0.610 6.449 0.264 1.216 8.935 1.615 1.788 6.443
@@ 4.1 Experimental Setup
To the best of our knowledge, no existing work addresses self-supervised articulated object modeling from few-shot unposed images. Nevertheless, we can evaluate our approach by comparing it with state-of-the-art methods that utilize posed images for articulated object modeling, specifically ArticulatedGS (AGS) [10]. For fair comparison, we employ VGGT [44] to estimate camera poses for the few-shot input images and point clouds initialization. We thus compare our method against two baselines: AGS with ground truth poses and AGS with VGGT estimated poses plus VGGT estimated point clouds, using 4 images per articulation state.
@@
4.2 Comparison with Baselines
Table 1 presents the quantitative comparison between our method and the baselines. The results clearly show that existing methods cannot handle the few-shot unposed setting—all three methods perform catastrophically across all metrics. Despite using VGGT-estimated point clouds for initialization, AGS-VGGT still fails to produce meaningful reconstructions. Our method, however, consistently outperforms the baselines across all metrics and object types. The baseline methods exhibit very high Chamfer Distance errors, indicating poor geometry reconstruction quality even at the source state. The poor performance stems from several fundamental limitations: limited input views severely hinder baseline reconstruction capabilities. Since these baseline methods heavily rely on precise 3D reconstruction and camera poses to identify articulated parts, the poor geometry reconstruction quality inevitably leads to inaccurate articulation parameter estimation and degraded novel articulation synthesis quality. As for the AYN method, it raises an error when running on the "Blade" object as it failed to initialize part segmentation due to poor reconstruction quality.
Despite outperforming the baselines, certain objects remain challenging for our method. For instance, the part segmentation of the "Blade" object shows limited accuracy, as illustrated in Fig. 5. We observe that only the blade tip and the blue handle are properly segmented, while the main blade body is classified as the root part. This occurs because the blade’s movement is too small, the pipeline could produce similar visual results during rendering by just moving the blade tip and the blue handle. Additionally, our method experiences performance degradation when dealing with thin articulated parts, like the lid of "Washer" and the door of "Oven". Since reconstructing thin objects is inherently challenging, the resulting geometry errors create larger margins for articulation estimation errors. Besides, objects with large occluded parts are difficult to to maintain consistency in the reconstructed geometry between source and target states. One example would be the USB, which has a large occluded part at target state. Such geometry inconsistency results in performance degradation.
We further examine the baseline failures through visual analysis. Figure 4 shows view synthesis results, where baseline methods generate heavily corrupted renderings with substantial noise—the objects become barely recognizable. Surprisingly, our method matches the quality of AGS-Full, which uses 100 images per articulation state with ground truth poses. Figure 5 displays the 3D Gaussian point distributions with part segmentation. The baselines produce scattered floating points and fragmented geometries, while our method creates coherent 3D structures with proper part segmentation and articulation parameter estimation.
@@ 4.2 Comparison with Baselines
rowsep=0pt,colsep=1pt RGB input AGS-GT AGS-VGGT AGS-Full Ours GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Visualization of 3D Gaussian points with part-level segmentation and articulation axes. Ground truth axises are shown in green, predictions in red. Better view in color and zoom in.
[…]
AGS GT 34.914 0.988 1.279 0.101
VGGT 28.860 0.975 1.367 1.623
AYN GT 32.219 0.962 0.946 0.654
Ours - 28.870 0.975 1.399 1.760
Table 2: Comparative analysis for 16 input views per articulation state.
Figure 6: PSNR w.r.t. number of views per articulation.
[…]
rowsep=1pt,colsep=1pt 4-view 8-view 16-view AGS-GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption AYN-GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption AGS-VGGT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Ours Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Qualitative evaluation for the effect of the number of input images. We show novel view synthesis for source state (left) and target state (right) at each settings. From top to bottom: AGS-GT, AYN-GT, AGS-VGGT, Ours.
[…]
rowsep=0pt,colsep=1pt $k_{0}=10^{-5}$ $k_{0}=10^{-8}$ $k_{0}=10^{-11}$ Refer to caption Refer to caption Refer to caption Coarse segmentation at source state Refer to caption Refer to caption Refer to caption Deformation Overlay of Gaussian Points
[…]
rowsep=0pt,colsep=1pt RGB input Novel articulation synthesis Part-level segmentation Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
4.3 Ablation Studies
@@
With just 4-view images, baseline methods struggle to reconstruct objects even when in static state. We investigate how the number of input views affects performance by varying the views per articulation state. Fig. 6 shows that PSNR for baseline methods improves with more input images, saturating after 16 images (VGGT runs out of memory with 64 images on our 48GB VRAM GPU), even with VGGT-estimated poses. Our method matches the saturated performance without ground truth poses with only 4 input images and stays consistent across different view counts with qualitative evaluation shows in Fig. 7. Tab. 2 shows the comparison at 16 images per state, where AGS with VGGT poses gets close to our performance. However, AGS still encounter reconstruction failures even with 32 images per state when VGGT provides less accurate poses. We suspect that AGS is more sensitive to pose estimation errors as it heavily depends on 2D image information. Our approach leverages FreeSplatter’s 3D priors and 3D-guided deformation, making it robust to pose estimation variations when poses stay consistent relative to the reconstructed splats.
Effect of $k_{0}$ for deformation field
4.4 Real-world Experiments
To demonstrate the generalization capability of our approach, we conduct experiments on real-world articulated objects. We evaluate our method on four different objects: a bottle, a projector, a drawer, and a lego toy, collecting four input images for each articulation state. Prior to processing, we employ the Segment Anything Model (SAM) [36] to remove background from the input images. For objects with multiple articulated parts, such as the lego toy with two movable arms, we collect three sets of input images per articulation state as explained in Sec. 3.4. The experimental results, presented in Fig. 9, demonstrate that our method successfully generates seamless reconstruction results for real-world articulated objects, validating its practical applicability beyond synthetic datasets.
5 Conclusion
We introduced PAOLI, a self-supervised framework for learning articulated object representations from sparse, unposed images. Unlike prior methods that require dense multi-view data and known camera poses, PAOLI operates with only a few views per articulation and no camera supervision. By combining sparse-view 3D reconstruction, a deformation field for cross-pose correspondences, and a progressive disentanglement strategy, our method jointly optimizes geometry, appearance, and kinematics while robustly separating camera and object motion.
Our work assumes articulated objects with rigid parts, cannot handle non-rigid deformations. Handling such transformations requires more complex deformation models. Second, we assume that the observations from two different poses of the same object are available. It would be valuable to pursue an active vision approach that allows the system to interact with the object while learning its articulated structure, which could improve robustness and generalization (e.g., [34]). Additionally, our current method relies on comprehensive object coverage with four input views per articulation state; insufficient coverage may degrade reconstruction quality and lead to failure. Integrating diffusion models to synthesize additional viewpoints could offer a promising solution for scenarios with limited view coverage. Finally, temporal information could be leveraged to improve the understanding of articulated motion dynamics.
Supplementary Material
6 Supplementary Material
6.1 Detailed Discussion of Related Work
Here we provide more detailed discussion of the most related articulated object learning work including, PARIS [22], AYN [7] and AGS [10]. As explained in the submission, these techniques assume dense views of the object across two articulation state along with the camera information, unlike our method which is pose-free and works with sparse views. Below, we explain the differences in methodology in more detail.
All the methods including ours uses the synthesis by analysis approach that deforms the object shape in the source articulation state to the target articulation state while learning the geometry, appearance and part segmentation as the intermediate representations. PARIS jointly optimizes the geometry, appearance and articulation parameters from the input images of both source and target states with photometric loss. It employs two separate NeRFs to model ‘static’ and ‘dynamic’ parts of the object. Although the joint optimization is straightforward, it is unstable and fails to obtain accurate results when the initialization of these parameters are noisy, as reported in both AYN and AGS. In addition, PARIS is limited to modeling one moving part only.
Instead, both AYN and AGS follows a multi-step approach, hence reconstructs the geometry and appearance of the object first in the source state. Once the initial geometry and appearance are learned, they use a decoupled optimization for part segmentation and motion parameter estimation that provides more stable optimization. AYN first extracts evenly spaced point cloud from NeRF scans, and then projects the point clouds back on image space. By comparing the difference in the alpha channel between the projected images and the target state images, AYN initializes the part segmentation label. With the initialized labels, it optimizes the part motion parameter so that the projected images with transformed part-points match the image of corresponding target images in alpha channel measured by 2D chamfer distance. Later, once the motion parameters are fixed, and AYN learns a segmentation field to refine the segmentation by minimizing the photo metric loss between rendered images and target images.
Similarly, AGS tackles the instability problem with multi-step approach. To first obtain the segmentation labels for gaussians, by employing a deformation field to learn the displacement for each Gaussian and rotation to correct the quaternions for color rendering. That’s $\hat{\bm{\mu}}=\bm{\mu}+\delta\bm{\mu}$, $\hat{\bm{q}}=\bm{q}\otimes\delta\bm{q}$. This steps aims to obtain the deformed 3D Gaussian that could render similar images in the target states with the corresponding camera poses. The deformation is supervised by photometric loss and a rigidity regularization term. Since the camera poses are aligned with the root part of the objects, the gaussian points in the root part will be static with small displacement $\delta\bm{\mu}$. AGS assigns Gaussians with small displacement to the static part, and the ones with large deformations to the moving part. After freezing their segmentation labels, it estimates a rigid transformation for the dynamic part with chamfer distance w.r.t. the deformed Gaussian (by deformation field) and photometric loss w.r.t. the images in the target states.
However, both AYN and AGS rely on known camera poses that are aligned at the root part between source and target states. This requirement prevents them from initializing part segmentation when camera poses are unknown across different articulation states.
Unlike previous methods that identify parts based on rendering differences (which only relies on 2D information from images) and assume a static root part, our method establishes correspondence across two articulation states directly in 3D space by deforming the 3D Gaussians from the source to the target state. Since $G^{s}$ and $\hat{G}^{t}$ represent the same set of Gaussians with $\hat{G}^{t}$ being the transformed version of $G^{s}$, we can naturally establish matching pairs between them. The deformation is supervised by 3D chamfer distance and photometric loss, ensuring that $\hat{G}^{t}$ closely approximates $G^{t}$. This makes the matching pairs from $G^{s}$ to $\hat{G}^{t}$ effectively equivalent to matching pairs between $G^{s}$ and $G^{t}$. With these correspondences established, we can separate parts and estimate per-part transformations using TEASER for initialization. To further improve rigidity during deformation and enhance performance, we employ positional encoding on the Gaussian means before feeding them to the deformation network.
With the initialization, we can now continue for fine-level optimization. For stability purpose, we also decouple the updates for part segmentation from the joint optimization. We split the optimization into 3 phases. The first phase we try to refine the coarse motion estimation from TEASER. In the second phase, we update part segmentation based on heuristic rules explained in §3.3 in the main paper to refine the segmentation. At the last phase, we optimize the appearance to improve the rendering quality.

### v2

@@ 3.4 Modeling Multiple Moving Parts
For simplicity, we have described our method for a two-part object. Our framework extends to objects with multiple parts by applying it iteratively, as shown in Fig. 3. This requires a common source image set ($\mathcal{I}^{s}$) and additional target sets ($\mathcal{I}^{t1},\mathcal{I}^{t2},\dots$), one for each moving part. We first run the full pipeline on the $(\mathcal{I}^{s},\mathcal{I}^{t1})$ pair to model the first part. We then repeat the process on the $(\mathcal{I}^{s},\mathcal{I}^{t2})$ pair to model the second part, and so on. The key assumption is that each source-target pair isolates the motion of only one part.
@@ 4 Experiments
Corr. Oven Storage USB Laptop
GaussReg Refer to caption Refer to caption Refer to caption Refer to caption
FM-Dense Refer to caption Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Qualitative analysis of correspondence methods. We visualize the part segmentations computed by our TEASER [59] solver, which is fed correspondences from each method. Inliers for the two main segmented parts are colored blue and orange; outliers and non-correspondences are shown in gray. Better view in color and zoom in.
@@ 4.1 Experimental Setup
GT AGS AGS-Full Ours
FM-Dense Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption failed Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption failed Refer to caption
@@
To our knowledge, no existing work addresses self-supervised articulated object modeling from sparse, unposed images. Our evaluation is therefore twofold: a component-wise analysis to validate our correspondence method and a full-pipeline comparison against state-of-the-art methods. For the component analysis, we fix our pipeline and swap our correspondence module with two alternatives: (1) GaussReg [3], a pre-trained rigid GS registration method that provides 1024 sparse keypoint correspondences, and (2) FM-Dense [2], a pre-trained non-rigid matching network, to which we feed 20480 downsampled points from our GS to obtain dense correspondences. For the full-pipeline comparison, we evaluate against ArticulatedGS (AGS) [16] with the same number of images ($K=4$) and ground truth, aligned camera parameters. Besides, we use the VGGT [53] reconstructed point clouds to initialize the AGS for better reconstruction. We will make our code and models publicly available.
4.2 Comparison with Baselines
Quantitative Evaluation
Quantitative results are presented in Tab. 1.
We first compare our method against ArticulatedGS (AGS) [16]. The results show that AGS, even with GT aligned camera poses and VGGT initialization, fails to converge in a sparse 4-view setting. This is because its optimization, designed for dense data, introduces significant noise from sparse inputs, and the limited views are insufficient for its motion estimation to converge. These factors lead to catastrophic errors, demonstrating that existing SOTA methods are not robust to this sparse-input scenario.
Our qualitative analysis in Fig. 4 visually confirms our quantitative findings by showing the output of the TEASER [59] solver for each baseline (inliers in blue/orange, outliers in gray). The GaussReg [3] output is entirely gray, confirming its 1024 sparse, noisy correspondences are insufficient to initialize segmentation. FM-Dense [2] proves highly unstable; the visualization shows most of its 20k correspondences are rejected as outliers, and even when a solution is found (e.g., Laptop), segmentation is incorrect, with inliers (blue/orange) assigned to the same part. In contrast, our method produces a dense, reliable map with a high percentage of correctly segmented inliers. A small volume of outliers is visible at part boundaries (e.g., USB, Storage), which are refined in the joint optimization step.
RGB input AGS-GT AGS-Full FM-Dense Ours GT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Visualization of 3D Gaussian points with part-level segmentation and articulation axes. Ground truth axises are shown in green, predictions in red. AGS-Full denotes the AGS pipeline with dense and posed images input. Better view in color and zoom in.
$k_{0}=10^{-5}$ $k_{0}=10^{-8}$ $k_{0}=10^{-11}$
Refer to caption Refer to caption Refer to caption
Coarse segmentation at source state
Refer to caption Refer to caption Refer to caption
Deformation Overlay of Gaussian Points
RGB input Novel articulation synthesis Part-level seg.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Qualitative Evaluation
We provide a qualitative analysis of the correspondence methods in Fig. 4. This figure visualizes the output of the TEASER [59] solver when fed correspondences from each baseline. Inliers for the two main parts are colored blue and orange, while outliers and non-correspondences are gray. The results visually confirm our quantitative findings. The GaussReg [3] baseline fails completely; its point cloud is rendered entirely gray, confirming that its 1024 sparse and noisy keypoint correspondences are insufficient to initialize segmentation. FM-Dense [2] proves highly unstable; the visualization shows most of its 20k correspondences are rejected as outliers, and even when a solution is found (e.g., Laptop), segmentation is incorrect, with inliers (blue/orange) assigned to the same part. In contrast, our method produces a dense, reliable map with a high percentage of correctly segmented inliers onto the distinct rigid parts. A small volume of outliers (gray points) is visible at the part boundaries on objects like the USB and Storage, which will can be further optimized by the joint optimization step.
Further qualitative evidence is provided in Fig. 6, which visualizes the final part segmentation and estimated joint axes. For this comparison, we include AGS [16] (using its full, dense-view and posed-image configuration) as a state-of-the-art upper bound. As the figure illustrates, our method’s joint parameter estimation is visually on par with the SOTA AGS baseline, despite our significantly more challenging sparse and unposed input. In contrast, the FM-Dense [2] baseline demonstrates degraded performance, resulting in clear errors in both segmentation and joint estimation. This visualization further highlights the robustness and accuracy of our full pipeline. We also provide the qualitative evaluation for novel articulation synthesis in Fig. 5. For more visualizations, please refer to the supplementary.
4.3 Ablation Studies
Effect of $k_{0}$ for deformation field
Real-world Experiments
To demonstrate the generalization of our approach, we conduct experiments on four real-world objects (a bottle, projector, drawer, and lego toy), capturing four input images per articulation state. We employ the Segment Anything Model (SAM) [43] to remove backgrounds prior to processing. For multi-part objects, like the lego toy, we capture three image sets (one per state) as detailed in Sec. 3.4. The results in Fig. 8 show that our method successfully generates seamless reconstructions for these real-world objects, validating its practical applicability.
5 Conclusion
We have introduced PAOLI, the first framework to successfully learn high-fidelity articulated object representations from sparse, unposed images. We identified that the central challenge in this setting is not just reconstruction, but the 3D-to-3D correspondence problem between unaligned 3D models.
Our solution is a test-time, optimization-based deformation field that finds dense and robust correspondences by leveraging both 3D geometry and 2D rendering losses. Our experiments show this approach is highly effective, whereas prominent baselines—including sparse keypoint methods and pre-trained dense networks—fail to solve the problem. Our full pipeline successfully recovers high-fidelity geometry, appearance, and kinematics, demonstrating a robust path forward for articulated object learning in practical, unposed scenarios.
While effective, our method assumes rigid parts and requires at least two distinct articulation states with adequate view coverage. Furthermore, our alignment may fail if the initial, arbitrary coordinate systems of the 3D GS reconstructions are extremely misaligned. More detailed discussion on limitations can be found in the supplementary.
[…]
Supplementary Material
6 Supplementary Material
6.1 Implementation Details
Deformation Network.
The deformation network $\mathcal{F}_{\text{deform}}$, introduced in Sec. 3.2, is a 3-layer MLP with 128-dimensional hidden layers and ReLU activations. The network outputs a rigid transformation where rotation is parameterized as Euler angles.
Computation Efficiency
Our full pipeline runs in approximately 15 minutes per object on a system with an AMD Zen3 7950X CPU and an NVIDIA RTX 4090 GPU. The runtime breaks down as follows:
  • Step 1: $\sim$1 minute for initial reconstruction and camera pose refinement.
  • Step 2: $\sim$4 minutes for deformation field optimization.
  • Step 3: $\sim$10 minutes, which includes a few seconds for the TEASER++ solver and the remainder for joint refinement.
The peak memory usage is approximately 15 GB of VRAM during the Step 3.
6.2 Ablation Studies
Full Pipeline w.o. Cam. Opt. w.o. $L_{CD}$ w.o. $L_{photo}$
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Visualization of the results after TEASER initialization in Step 3. The main impact of these ablated components occurs before this initialization step. Better view in color and zoom in.
Impact of Key Pipeline Components.
Table 2: Ablation study on the key components of our pipeline. We report the average metrics across all objects. Removing any component results in performance degradation, highlighting their importance.
Components Metrics
cam_opt $\mathcal{L}_{\text{CD}}$ $\mathcal{L}_{\text{photo}}$ Ang Err $\downarrow$ Pos Err $\downarrow$ Motion Dist $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ CD $\downarrow$
$\checkmark$ $\checkmark$ $\checkmark$ 1.160 0.105 1.318 26.573 0.953 2.272
@@ 6.2 Ablation Studies
We conducted an ablation study to analyze the impact of key components in our pipeline: camera pose refinement in step 1 and the losses used for deformation network optimization ($\mathcal{L}_{\text{CD}}$ and $\mathcal{L}_{\text{photo}}$) in step 2. As shown in Tab. 2, removing any of these components leads to a degradation in performance.
These components are crucial for initializing the part segmentation and pose estimation in Step 3. To understand the performance drop, we examined the initialization results. As illustrated in Fig. 9, omitting these components leads to poorer initializations, which subsequently hampers the final joint optimization stage. This demonstrates that each component plays a vital role in the robustness of our framework.
Effectiveness of the Joint Optimization
The effectiveness of our joint optimization process is demonstrated in Fig. 10. The figure shows how the process corrects initial errors: despite dense matching, the state after TEASER initialization still has spatial discrepancies between parts. Phase 1 recovers connectivity, Phase 2 refines part segmentation, and Phase 3 produces the final, rigid, joint-optimized output.
TEASER init. Phase 1 Phase 2 Phase 3
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Visualization of the effectiveness of the joint optimization process, showing the progression from the initial state to the final refined output.
@@
With just 4-view images, traditional articulated object modelling methods struggle to reconstruct objects even when in static state. We investigate how the number of input views affects performance by varying the views per articulation state. Fig. 11 shows the rendering quality for ArticulatedGS [16] methods improves with more input images, saturating after 16 images (VGGT runs out of memory with 64 images on our 48GB VRAM GPU), even with VGGT-estimated poses. Our method matches the saturated performance without ground truth poses with only 4 input images and stays consistent across different view counts with qualitative evaluation.
More Qualitative Results
We provide additional qualitative results for correspondences and segmentation in Fig. 13 and Fig. 15, respectively. Besides, correspondences visualization on the real world images are also shown in Fig. 14.
6.3 Failure Cases
Extreme Initial Misalignment.
We demonstrate some failure cases in Fig. 12. Our method relies on FreeSplatter [62] for initial 3D reconstruction, which establishes a coordinate system based on the first input image of a set. If the first images from the source and target sets are captured from drastically different viewpoints (e.g., front vs. back), the resulting 3D models can be severely misaligned. This extreme misalignment can cause our deformation network to converge to a poor local minimum, especially for objects with geometric symmetries. For instance, as shown with the real-world storage drawer in Fig. 7, providing opposing side views as initial frames causes the network to favor a non-rigid deformation of the drawer rather than finding the correct global alignment. To mitigate this, we recommend selecting initial views from similar perspectives for each image set.
4-view 8-view 16-view
AGS-GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
AGS-VGGT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Qualitative evaluation for the effect of the number of input images. We show novel view synthesis for source state (left) and target state (right) at each settings. From top to bottom: AGS-GT, AGS-VGGT, Ours.
Input reference view Render deformation target states 3D overlay
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: Visualization of the failure cases in extreme misalignment case (top) and thin structures with large motion (bottom). The 3D overlay shows the itermediate state during the optimization of deformation field, the red is the deformed source state, the green is the target state.
Thin Structures with Large Motion.
The framework can struggle with objects that have very thin parts undergoing large articulation. A thin structure is represented by a smaller volume of Gaussian splats, which provides weaker rigidity constraints during the deformation step. When combined with a large motion, the deformation field may fail to preserve the part’s rigidity, leading to inaccurate correspondence and subsequent errors in joint estimation. An example of this can be seen in Fig. 12.